OpenStudy (anonymous):

you are going to a 4 year college in 4 years that will cost $14,895.00 yr. your parents expect you to pay 5% of the total cost. 3. how much do you need to pay each year of attending? 4. if you want to save total contribution for all 4 years befor you start attending college, how much do you need to save each month if you have 4 years to accomplish your goal?

OpenStudy (espex):

What part are you stuck on?

OpenStudy (anonymous):

all of it im not good in math

OpenStudy (espex):

First, you are being asked to pay 5% of the total cost. So what is 5% of 14895.00?

OpenStudy (anonymous):

74475 is it this

OpenStudy (anonymous):

@eSpeX

OpenStudy (espex):

Adjust your decimal place, but yes, that is 5%

OpenStudy (espex):

So question 3 is then asking you how much would you have to pay each of the 4 years to have it paid off by the time you graduate.

OpenStudy (espex):

So what is your 5% divided by 4?

OpenStudy (anonymous):

20

OpenStudy (espex):

heh, no, your 5%, which equals 744.75. :)

OpenStudy (anonymous):

3723.75

OpenStudy (espex):

How did you get that?

OpenStudy (anonymous):

by multiplying

OpenStudy (espex):

You are asked, what is 5% of 14895.00, in math speak that is x= 0.05 * 14895.00 x = 744.75 Then you want to know how much you would pay each of the 4 years to get that 744.75. In math speak you are asked, x = 744.75/4 So what is x?

OpenStudy (anonymous):

186.1875

OpenStudy (espex):

Exactly.

OpenStudy (espex):

Looking at the problem statement, it looks like it says 14,895.00 per year, is that right? It will change the numbers a little but we can use the same process.

OpenStudy (anonymous):

yea

OpenStudy (espex):

Okay, so it will change slightly then.

OpenStudy (espex):

To start you want ot find the total tuition. 14,895.55 * 4 = 59,580.00 Then 5% of 59598, 0.05*59580= 2979.00 2979/4 = 744.75 a year must be paid.

OpenStudy (espex):

Then question 4 asks you how much each month would you need to save if you were paying the total contribution before you were to attend.

OpenStudy (espex):

Since you have 4 years to save, and you are asked about a monthly sum, how many months are there in 4 years?

OpenStudy (anonymous):

wait im lost so #3 is 186.1875 ?

OpenStudy (espex):

No, that was my bad as I misread the question. I read a total tuition of 14895, when in fact that was a yearly tuition.

OpenStudy (espex):

Since they gave you 14,895.00 per year, then your total tuition was 59,580.00 and you are asked to pay 5% of that. 59580 * 0.05 = 2979.00 and dividing that by 4 gives you 744.75 per year. So the answer to 3 is $744.75 per year.

OpenStudy (anonymous):

ok thank you I was lost so number for is wat he save know ?

OpenStudy (espex):

Number 4 is asking you how much you need to save each month to get your total contribution.

OpenStudy (espex):

You have 4 years to save 2979.00, how much each month will you need to save?

OpenStudy (anonymous):

ok so I just subtuct then right ?

OpenStudy (espex):

First, how many months are there in 4 years?

OpenStudy (anonymous):

48

OpenStudy (espex):

Good, now you only need to divide your 2979 over 48 months to get your answer. In other words, what is 2979/48?